OpenStudy (anonymous):

find all zeros of the function algebraically 2x^2-7x-30

OpenStudy (anonymous):

Find when \(2x^2-7x-30=0\). There are a few ways to do this. What methods are you familiar with?

OpenStudy (anonymous):

any method is fine

OpenStudy (anonymous):

splitting would be the best method, on go on substituting with possible zeroes

OpenStudy (anonymous):

how would you split it?

OpenStudy (anonymous):

right, v cant, i just found out!!!!!!!!!!

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

you can!! 2x^2-12x+5x-30 = 0.............. here!!

OpenStudy (anonymous):

i hope u gt it, i cud explain!

OpenStudy (anonymous):

could you please...

OpenStudy (anonymous):

sure.... the product is 60- ( a x c = 12 x 5) and 12 - 5 is also 7, so you split the 7x as -12x+5x, and finally you can find by taking common factors out!